# atomic orbital API Application Programmer Interface

### v1

@@ 4.1. Algorithm
$\delta\varepsilon_{11}\varphi_{1}+\delta\varepsilon_{12}\varphi_{2}+\varepsilon_{11}\delta\varphi_{1}+\varepsilon_{12}\delta\varphi_{2}=F(\delta\varphi_{1},\delta\varphi_{2},\delta\theta,w)$ (4.3)
$\delta\varepsilon_{12}\varphi_{1}+\delta\varepsilon_{22}\varphi_{2}+\varepsilon_{12}\delta\varphi_{1}+\varepsilon_{22}\delta\varphi_{2}=G(\delta\varphi_{1},\delta\varphi_{2},\delta\theta,w)$
@@
$\begin{pmatrix}\left\|\varphi_{1}\right\|^{2}&0&\int\varphi_{1}\varphi_{2}&0&\varepsilon_{12}\\ 0&\left\|\varphi_{2}\right\|^{2}&\int\varphi_{1}\varphi_{2}&\varepsilon_{12}&0\\ \int\varphi_{1}\varphi_{2}&0&\left\|\varphi_{2}\right\|^{2}&\varepsilon_{11}&0\\ 0&\int\varphi_{1}\varphi_{2}&\left\|\varphi_{1}\right\|^{2}&0&\varepsilon_{22}\\ 0&0&0&1&1\end{pmatrix}\begin{pmatrix}\delta\varepsilon_{11}\\ \delta\varepsilon_{22}\\ \delta\varepsilon_{12}\\ \int\delta\varphi_{1}\varphi_{2}\\ \int\varphi_{1}\delta\varphi_{2}\end{pmatrix}=\begin{pmatrix}\int F\varphi_{1}-\frac{1}{2}\varepsilon_{11}\left(1-\left\|\varphi_{1}\right\|^{2}\right)\\ \int G\varphi_{2}-\frac{1}{2}\varepsilon_{22}\left(1-\left\|\varphi_{2}\right\|^{2}\right)\\ \int F\varphi_{2}-\frac{1}{2}\varepsilon_{12}\left(1-\left\|\varphi_{2}\right\|^{2}\right)\\ \int G\varphi_{1}-\frac{1}{2}\varepsilon_{12}\left(1-\left\|\varphi_{1}\right\|^{2}\right)\\ -\int\varphi_{1}\varphi_{2}\end{pmatrix}$ (4.4)
@@
We search for orthonormal orbitals $\varphi_{1},\varphi_{2}$. Therefore, the determinant of this system should stay close to $-\varepsilon_{11}-\varepsilon_{22}$ during each Newton’s update. In particular, the matrix should be invertible for each Newton’s iteration $w=w_{n}$.
$\int F\varphi_{1}=c_{1}^{2}(\delta 1|h|1)+3c_{1}^{2}(\delta 11|11)+c_{1}c_{2}(\delta 12|12)+2c_{1}c_{2}(1\delta 2|12)\\ +(c_{1}^{2}-\delta\theta\sin 2\theta)((1|h|1)+(11|11))+(c_{1}c_{2}+\delta\theta\cos 2\theta)(12|12)-\varepsilon_{11}\left\|\varphi_{1}\right\|^{2}-\varepsilon_{12}\int\varphi_{1}\varphi_{2}$ (4.5)
$\int F\varphi_{2}=c_{1}^{2}(\delta 1|h|2)+c_{1}^{2}(\delta 12|11)+2c_{1}^{2}(\delta 11|12)+c_{1}c_{2}(\delta 12|22)+c_{1}c_{2}(1\delta 2|22)+c_{1}c_{2}(2\delta 2|12)\\ +(c_{1}^{2}-\delta\theta\sin 2\theta)((1|h|2)+(11|12))+(c_{1}c_{2}+\delta\theta\cos 2\theta)(12|22)-\varepsilon_{11}\int\varphi_{1}\varphi_{2}-\varepsilon_{12}\left\|\varphi_{2}\right\|^{2}$ (4.6)
$\int G\varphi_{2}=c_{2}^{2}(\delta 2|h|2)+3c_{2}^{2}(\delta 22|22)+2c_{1}c_{2}(\delta 12|12)+c_{1}c_{2}(1\delta 2|12)\\ +(c_{2}^{2}+\delta\theta\sin 2\theta)((2|h|2)+(22|22))+(c_{1}c_{2}+\delta\theta\cos 2\theta)(12|12)-\varepsilon_{22}\left\|\varphi_{2}\right\|^{2}-\varepsilon_{12}\int\varphi_{1}\varphi_{2}$ (4.7)
$\int G\varphi_{1}=c_{2}^{2}(\delta 2|h|1)+c_{2}^{2}(1\delta 2|22)+2c_{2}^{2}(\delta 22|12)+c_{1}c_{2}(\delta 12|11)+c_{1}c_{2}(1\delta 2|11)+c_{1}c_{2}(1\delta 1|12)\\ +(c_{2}^{2}+\delta\theta\sin 2\theta)((1|h|2)+(12|22))+(c_{1}c_{2}+\delta\theta\cos 2\theta)(11|12)-\varepsilon_{22}\int\varphi_{1}\varphi_{2}-\varepsilon_{12}\left\|\varphi_{1}\right\|^{2}$ (4.8)
$c_{1}^{2}\left(h+\frac{1}{|x|}*\varphi_{1}^{2}\right)\delta\varphi_{1}+2c_{1}^{2}\left(\frac{1}{|x|}*(\delta\varphi_{1}\varphi_{1})\right)\varphi_{1}+c_{1}c_{2}\left(\frac{1}{|x|}*(\delta\varphi_{1}\varphi_{2})\right)\varphi_{2}-\varepsilon_{11}\delta\varphi_{1}\\ +c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\delta\varphi_{2})\right)\varphi_{2}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\delta\varphi_{2}-\varepsilon_{12}\delta\varphi_{2}\\ -\delta\theta\sin 2\theta\left(h+\frac{1}{|x|}*\varphi_{1}^{2}\right)\varphi_{1}+\delta\theta\cos 2\theta\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{2}-\delta\varepsilon_{11}\varphi_{1}-\delta\varepsilon_{12}\varphi_{2}\\ +c_{1}^{2}\left(h+\frac{1}{|x|}*\varphi_{1}^{2}\right)\varphi_{1}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{2}-\varepsilon_{11}\varphi_{1}-\varepsilon_{12}\varphi_{2}=0$ (4.9)
$c_{1}c_{2}\left(\frac{1}{|x|}*(\delta\varphi_{1}\varphi_{2})\right)\varphi_{1}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\delta\varphi_{1}-\varepsilon_{12}\delta\varphi_{1}\\ +c_{2}^{2}\left(h+\frac{1}{|x|}*\varphi_{2}^{2}\right)\delta\varphi_{2}+2c_{2}^{2}\left(\frac{1}{|x|}*(\delta\varphi_{2}\varphi_{2})\right)\varphi_{2}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\delta\varphi_{2})\right)\varphi_{1}-\varepsilon_{22}\delta\varphi_{2}\\ +\delta\theta\sin 2\theta\left(h+\frac{1}{|x|}*\varphi_{2}^{2}\right)\varphi_{2}+\delta\theta\cos 2\theta\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{1}-\delta\varepsilon_{12}\varphi_{1}-\delta\varepsilon_{22}\varphi_{2}\\ +c_{2}^{2}\left(h+\frac{1}{|x|}*\varphi_{2}^{2}\right)\varphi_{2}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{1}-\varepsilon_{12}\varphi_{1}-\varepsilon_{22}\varphi_{2}=0$ (4.10)
$c_{1}^{2}\left(h+J(11)\right)\delta\varphi_{1}+2c_{1}^{2}J(1\delta 1)\varphi_{1}+c_{1}c_{2}J(\delta 12)\varphi_{2}-\varepsilon_{11}\delta\varphi_{1}\\ +c_{1}c_{2}J(1\delta 2)\varphi_{2}+c_{1}c_{2}J(12)\delta\varphi_{2}-\varepsilon_{12}\delta\varphi_{2}\\ -\delta\theta\sin 2\theta\left(h+J(11)\right)\varphi_{1}+\delta\theta\cos 2\theta J(12)\varphi_{2}-\delta\varepsilon_{11}\varphi_{1}-\delta\varepsilon_{12}\varphi_{2}\\ +c_{1}^{2}\left(h+J(11)\right)\varphi_{1}+c_{1}c_{2}J(12)\varphi_{2}-\varepsilon_{11}\varphi_{1}-\varepsilon_{12}\varphi_{2}=0$ (4.11)
$c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(12)\delta\varphi_{1}-\varepsilon_{12}\delta\varphi_{1}\\ +c_{2}^{2}\left(h+J(22)\right)\delta\varphi_{2}+2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{1}-\varepsilon_{22}\delta\varphi_{2}\\ +\delta\theta\sin 2\theta\left(h+J(22)\right)\varphi_{2}+\delta\theta\cos 2\theta J(12)\varphi_{1}-\delta\varepsilon_{12}\varphi_{1}-\delta\varepsilon_{22}\varphi_{2}\\ +c_{2}^{2}\left(h+J(22)\right)\varphi_{2}+c_{1}c_{2}J(12)\varphi_{1}-\varepsilon_{12}\varphi_{1}-\varepsilon_{22}\varphi_{2}=0$ (4.12)
@@
$c_{1}^{2}\left(-\frac{1}{2}\Delta+V_{\text{nuc}}+J(11)\right)\delta\varphi_{1}+2c_{1}^{2}J(1\delta 1)\varphi_{1}+c_{1}c_{2}J(\delta 12)\varphi_{2}-\varepsilon_{11}\delta\varphi_{1}\\ +c_{1}c_{2}J(1\delta 2)\varphi_{2}+c_{1}c_{2}J(12)\delta\varphi_{2}-\varepsilon_{12}\delta\varphi_{2}\\ -\delta\varepsilon_{11}\varphi_{1}-\delta\varepsilon_{12}\varphi_{2}-\varepsilon_{11}\varphi_{1}-\varepsilon_{12}\varphi_{2}\\ +\left(c_{1}^{2}-\delta\theta\sin 2\theta\right)\left(-\frac{1}{2}\Delta+V_{\text{nuc}}+J(11)\right)\varphi_{1}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{2}=0$ (4.13)
@@
$c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(12)\delta\varphi_{1}-\varepsilon_{12}\delta\varphi_{1}\\ +c_{2}^{2}\left(-\frac{1}{2}\Delta+V_{\text{nuc}}+J(22)\right)\delta\varphi_{2}+2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{1}-\varepsilon_{22}\delta\varphi_{2}\\ -\delta\varepsilon_{12}\varphi_{1}-\delta\varepsilon_{22}\varphi_{2}-\varepsilon_{12}\varphi_{1}-\varepsilon_{22}\varphi_{2}\\ +\left(c_{2}^{2}+\delta\theta\sin 2\theta\right)\left(-\frac{1}{2}\Delta+V_{\text{nuc}}+J(22)\right)\varphi_{2}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{1}=0$ (4.14)
@@
Finally, we rewrite these two equations in the self consistent form
@@
$\left(-\frac{c_{1}^{2}}{2}\Delta-\varepsilon_{11}\right)\delta\varphi_{1}+c_{1}^{2}\left(V_{\text{nuc}}+J(11)\right)\delta\varphi_{1}\\ +2c_{1}^{2}J(1\delta 1)\varphi_{1}+c_{1}c_{2}J(\delta 12)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{2}+c_{1}c_{2}J(12)\delta\varphi_{2}-\varepsilon_{12}\delta\varphi_{2}\\ -\delta\varepsilon_{11}\varphi_{1}-\delta\varepsilon_{12}\varphi_{2}-\varepsilon_{11}\varphi_{1}-\varepsilon_{12}\varphi_{2}\\ +\left(1-\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}\right)\left(-\frac{c_{1}^{2}}{2}\Delta\right)\varphi_{1}+\left(c_{1}^{2}-\delta\theta\sin 2\theta\right)\left(V_{\text{nuc}}+J(11)\right)\varphi_{1}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{2}=0$ (4.15)
@@
$c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(12)\delta\varphi_{1}+2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{1}-\varepsilon_{12}\delta\varphi_{1}\\ +\left(-\frac{c_{2}^{2}}{2}\Delta-\varepsilon_{22}\right)\delta\varphi_{2}+c_{2}^{2}\left(V_{\text{nuc}}+J(22)\right)\delta\varphi_{2}\\ -\delta\varepsilon_{12}\varphi_{1}-\delta\varepsilon_{22}\varphi_{2}-\varepsilon_{12}\varphi_{1}-\varepsilon_{22}\varphi_{2}\\ +\left(1+\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}\right)\left(-\frac{c_{2}^{2}}{2}\Delta\right)\varphi_{2}+\left(c_{2}^{2}+\delta\theta\sin 2\theta\right)\left(V_{\text{nuc}}+J(22)\right)\varphi_{2}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{1}=0$ (4.16)
Thus
@@
$\left(-\frac{c_{1}^{2}}{2}\Delta-\varepsilon_{11}\right)\delta\varphi_{1}+\left(1-\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}\right)\left(-\frac{c_{1}^{2}}{2}\Delta-\varepsilon_{11}\right)\varphi_{1}\\ +c_{1}^{2}\left(V_{\text{nuc}}+J(11)\right)\delta\varphi_{1}+2c_{1}^{2}J(1\delta 1)\varphi_{1}+c_{1}c_{2}J(\delta 12)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{2}+c_{1}c_{2}J(12)\delta\varphi_{2}-\varepsilon_{12}\delta\varphi_{2}\\ -\delta\varepsilon_{12}\varphi_{2}-\varepsilon_{12}\varphi_{2}-\left(\varepsilon_{11}\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}+\delta\varepsilon_{11}\right)\varphi_{1}\\ +\left(c_{1}^{2}-\delta\theta\sin 2\theta\right)\left(V_{\text{nuc}}+J(11)\right)\varphi_{1}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{2}=0$ (4.17)

### v2

@@ 3.1. Self-consistent form
(3.4) $\delta\varepsilon_{11}\varphi_{1}+\delta\varepsilon_{12}\varphi_{2}+\varepsilon_{11}\delta\varphi_{1}+\varepsilon_{12}\delta\varphi_{2}=F(\delta\varphi_{1},\delta\varphi_{2},\delta\theta,w),$
$\delta\varepsilon_{12}\varphi_{1}+\delta\varepsilon_{22}\varphi_{2}+\varepsilon_{12}\delta\varphi_{1}+\varepsilon_{22}\delta\varphi_{2}=G(\delta\varphi_{1},\delta\varphi_{2},\delta\theta,w).$
@@
(3.5) $\begin{pmatrix}\left\|\varphi_{1}\right\|^{2}&0&\int\varphi_{1}\varphi_{2}&0&\varepsilon_{12}\\ 0&\left\|\varphi_{2}\right\|^{2}&\int\varphi_{1}\varphi_{2}&\varepsilon_{12}&0\\ \int\varphi_{1}\varphi_{2}&0&\left\|\varphi_{2}\right\|^{2}&\varepsilon_{11}&0\\ 0&\int\varphi_{1}\varphi_{2}&\left\|\varphi_{1}\right\|^{2}&0&\varepsilon_{22}\\ 0&0&0&1&1\end{pmatrix}\begin{pmatrix}\delta\varepsilon_{11}\\ \delta\varepsilon_{22}\\ \delta\varepsilon_{12}\\ \int\delta\varphi_{1}\varphi_{2}\\ \int\varphi_{1}\delta\varphi_{2}\end{pmatrix}=\begin{pmatrix}\int F\varphi_{1}-\frac{1}{2}\varepsilon_{11}\left(1-\left\|\varphi_{1}\right\|^{2}\right)\\ \int G\varphi_{2}-\frac{1}{2}\varepsilon_{22}\left(1-\left\|\varphi_{2}\right\|^{2}\right)\\ \int F\varphi_{2}-\frac{1}{2}\varepsilon_{12}\left(1-\left\|\varphi_{2}\right\|^{2}\right)\\ \int G\varphi_{1}-\frac{1}{2}\varepsilon_{12}\left(1-\left\|\varphi_{1}\right\|^{2}\right)\\ -\int\varphi_{1}\varphi_{2}\end{pmatrix}$
@@
We search for orthonormal orbitals $\varphi_{1},\varphi_{2}$. Therefore, the determinant of this system should stay close to $\varepsilon_{11}-\varepsilon_{22}$ during each Newton’s update. In particular, it ensures that the matrix is invertible for each Newton’s iteration $w=w_{n}$. The integrals staying in the right hand side of (3.5) are computed as follows
(3.6) $\int F\varphi_{1}=c_{1}^{2}(\delta 1|h|1)+3c_{1}^{2}(\delta 11|11)+c_{1}c_{2}(\delta 12|12)+2c_{1}c_{2}(1\delta 2|12)\\ +(c_{1}^{2}-\delta\theta\sin 2\theta)((1|h|1)+(11|11))+(c_{1}c_{2}+\delta\theta\cos 2\theta)(12|12)-\varepsilon_{11}\left\|\varphi_{1}\right\|^{2}-\varepsilon_{12}\int\varphi_{1}\varphi_{2},$
(3.7) $\int F\varphi_{2}=c_{1}^{2}(\delta 1|h|2)+c_{1}^{2}(\delta 12|11)+2c_{1}^{2}(\delta 11|12)+c_{1}c_{2}(\delta 12|22)+c_{1}c_{2}(1\delta 2|22)+c_{1}c_{2}(2\delta 2|12)\\ +(c_{1}^{2}-\delta\theta\sin 2\theta)((1|h|2)+(11|12))+(c_{1}c_{2}+\delta\theta\cos 2\theta)(12|22)-\varepsilon_{11}\int\varphi_{1}\varphi_{2}-\varepsilon_{12}\left\|\varphi_{2}\right\|^{2},$
(3.8) $\int G\varphi_{2}=c_{2}^{2}(\delta 2|h|2)+3c_{2}^{2}(\delta 22|22)+2c_{1}c_{2}(\delta 12|12)+c_{1}c_{2}(1\delta 2|12)\\ +(c_{2}^{2}+\delta\theta\sin 2\theta)((2|h|2)+(22|22))+(c_{1}c_{2}+\delta\theta\cos 2\theta)(12|12)-\varepsilon_{22}\left\|\varphi_{2}\right\|^{2}-\varepsilon_{12}\int\varphi_{1}\varphi_{2},$
(3.9) $\int G\varphi_{1}=c_{2}^{2}(\delta 2|h|1)+c_{2}^{2}(1\delta 2|22)+2c_{2}^{2}(\delta 22|12)+c_{1}c_{2}(\delta 12|11)+c_{1}c_{2}(1\delta 2|11)+c_{1}c_{2}(1\delta 1|12)\\ +(c_{2}^{2}+\delta\theta\sin 2\theta)((1|h|2)+(12|22))+(c_{1}c_{2}+\delta\theta\cos 2\theta)(11|12)-\varepsilon_{22}\int\varphi_{1}\varphi_{2}-\varepsilon_{12}\left\|\varphi_{1}\right\|^{2}.$
It is left to rearrange the functional equations
(3.10) $c_{1}^{2}\left(h+\frac{1}{|x|}*\varphi_{1}^{2}\right)\delta\varphi_{1}+2c_{1}^{2}\left(\frac{1}{|x|}*(\delta\varphi_{1}\varphi_{1})\right)\varphi_{1}+c_{1}c_{2}\left(\frac{1}{|x|}*(\delta\varphi_{1}\varphi_{2})\right)\varphi_{2}-\varepsilon_{11}\delta\varphi_{1}\\ +c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\delta\varphi_{2})\right)\varphi_{2}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\delta\varphi_{2}-\varepsilon_{12}\delta\varphi_{2}\\ -\delta\theta\sin 2\theta\left(h+\frac{1}{|x|}*\varphi_{1}^{2}\right)\varphi_{1}+\delta\theta\cos 2\theta\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{2}-\delta\varepsilon_{11}\varphi_{1}-\delta\varepsilon_{12}\varphi_{2}\\ +c_{1}^{2}\left(h+\frac{1}{|x|}*\varphi_{1}^{2}\right)\varphi_{1}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{2}-\varepsilon_{11}\varphi_{1}-\varepsilon_{12}\varphi_{2}=0$
and
(3.11) $c_{1}c_{2}\left(\frac{1}{|x|}*(\delta\varphi_{1}\varphi_{2})\right)\varphi_{1}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\delta\varphi_{1}-\varepsilon_{12}\delta\varphi_{1}\\ +c_{2}^{2}\left(h+\frac{1}{|x|}*\varphi_{2}^{2}\right)\delta\varphi_{2}+2c_{2}^{2}\left(\frac{1}{|x|}*(\delta\varphi_{2}\varphi_{2})\right)\varphi_{2}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\delta\varphi_{2})\right)\varphi_{1}-\varepsilon_{22}\delta\varphi_{2}\\ +\delta\theta\sin 2\theta\left(h+\frac{1}{|x|}*\varphi_{2}^{2}\right)\varphi_{2}+\delta\theta\cos 2\theta\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{1}-\delta\varepsilon_{12}\varphi_{1}-\delta\varepsilon_{22}\varphi_{2}\\ +c_{2}^{2}\left(h+\frac{1}{|x|}*\varphi_{2}^{2}\right)\varphi_{2}+c_{1}c_{2}\left(\frac{1}{|x|}*(\varphi_{1}\varphi_{2})\right)\varphi_{1}-\varepsilon_{12}\varphi_{1}-\varepsilon_{22}\varphi_{2}=0.$
Using the short notation $J$ for convolutions, these two equations can rewritten as
(3.12) $c_{1}^{2}\left(h+J(11)\right)\delta\varphi_{1}+2c_{1}^{2}J(1\delta 1)\varphi_{1}+c_{1}c_{2}J(\delta 12)\varphi_{2}-\varepsilon_{11}\delta\varphi_{1}\\ +c_{1}c_{2}J(1\delta 2)\varphi_{2}+c_{1}c_{2}J(12)\delta\varphi_{2}-\varepsilon_{12}\delta\varphi_{2}\\ -\delta\theta\sin 2\theta\left(h+J(11)\right)\varphi_{1}+\delta\theta\cos 2\theta J(12)\varphi_{2}-\delta\varepsilon_{11}\varphi_{1}-\delta\varepsilon_{12}\varphi_{2}\\ +c_{1}^{2}\left(h+J(11)\right)\varphi_{1}+c_{1}c_{2}J(12)\varphi_{2}-\varepsilon_{11}\varphi_{1}-\varepsilon_{12}\varphi_{2}=0$
and
(3.13) $c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(12)\delta\varphi_{1}-\varepsilon_{12}\delta\varphi_{1}\\ +c_{2}^{2}\left(h+J(22)\right)\delta\varphi_{2}+2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{1}-\varepsilon_{22}\delta\varphi_{2}\\ +\delta\theta\sin 2\theta\left(h+J(22)\right)\varphi_{2}+\delta\theta\cos 2\theta J(12)\varphi_{1}-\delta\varepsilon_{12}\varphi_{1}-\delta\varepsilon_{22}\varphi_{2}\\ +c_{2}^{2}\left(h+J(22)\right)\varphi_{2}+c_{1}c_{2}J(12)\varphi_{1}-\varepsilon_{12}\varphi_{1}-\varepsilon_{22}\varphi_{2}=0.$
Equivalently,
@@
(3.14) $c_{1}^{2}\left(-\frac{1}{2}\Delta+V_{\text{nuc}}+J(11)\right)\delta\varphi_{1}+2c_{1}^{2}J(1\delta 1)\varphi_{1}+c_{1}c_{2}J(\delta 12)\varphi_{2}-\varepsilon_{11}\delta\varphi_{1}\\ +c_{1}c_{2}J(1\delta 2)\varphi_{2}+c_{1}c_{2}J(12)\delta\varphi_{2}-\varepsilon_{12}\delta\varphi_{2}\\ -\delta\varepsilon_{11}\varphi_{1}-\delta\varepsilon_{12}\varphi_{2}-\varepsilon_{11}\varphi_{1}-\varepsilon_{12}\varphi_{2}\\ +\left(c_{1}^{2}-\delta\theta\sin 2\theta\right)\left(-\frac{1}{2}\Delta+V_{\text{nuc}}+J(11)\right)\varphi_{1}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{2}=0$
and
@@
(3.15) $c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(12)\delta\varphi_{1}-\varepsilon_{12}\delta\varphi_{1}\\ +c_{2}^{2}\left(-\frac{1}{2}\Delta+V_{\text{nuc}}+J(22)\right)\delta\varphi_{2}+2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{1}-\varepsilon_{22}\delta\varphi_{2}\\ -\delta\varepsilon_{12}\varphi_{1}-\delta\varepsilon_{22}\varphi_{2}-\varepsilon_{12}\varphi_{1}-\varepsilon_{22}\varphi_{2}\\ +\left(c_{2}^{2}+\delta\theta\sin 2\theta\right)\left(-\frac{1}{2}\Delta+V_{\text{nuc}}+J(22)\right)\varphi_{2}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{1}=0.$
@@
Finally, we rewrite these two equations in the self consistent form, by first regrupping the terms as follows
@@
(3.16) $\left(-\frac{c_{1}^{2}}{2}\Delta-\varepsilon_{11}\right)\delta\varphi_{1}+c_{1}^{2}\left(V_{\text{nuc}}+J(11)\right)\delta\varphi_{1}\\ +2c_{1}^{2}J(1\delta 1)\varphi_{1}+c_{1}c_{2}J(\delta 12)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{2}+c_{1}c_{2}J(12)\delta\varphi_{2}-\varepsilon_{12}\delta\varphi_{2}\\ -\delta\varepsilon_{11}\varphi_{1}-\delta\varepsilon_{12}\varphi_{2}-\varepsilon_{11}\varphi_{1}-\varepsilon_{12}\varphi_{2}\\ +\left(1-\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}\right)\left(-\frac{c_{1}^{2}}{2}\Delta\right)\varphi_{1}+\left(c_{1}^{2}-\delta\theta\sin 2\theta\right)\left(V_{\text{nuc}}+J(11)\right)\varphi_{1}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{2}=0,$
@@
(3.17) $c_{1}c_{2}J(\delta 12)\varphi_{1}+c_{1}c_{2}J(12)\delta\varphi_{1}+2c_{2}^{2}J(2\delta 2)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{1}-\varepsilon_{12}\delta\varphi_{1}\\ +\left(-\frac{c_{2}^{2}}{2}\Delta-\varepsilon_{22}\right)\delta\varphi_{2}+c_{2}^{2}\left(V_{\text{nuc}}+J(22)\right)\delta\varphi_{2}\\ -\delta\varepsilon_{12}\varphi_{1}-\delta\varepsilon_{22}\varphi_{2}-\varepsilon_{12}\varphi_{1}-\varepsilon_{22}\varphi_{2}\\ +\left(1+\frac{\delta\theta\sin 2\theta}{c_{2}^{2}}\right)\left(-\frac{c_{2}^{2}}{2}\Delta\right)\varphi_{2}+\left(c_{2}^{2}+\delta\theta\sin 2\theta\right)\left(V_{\text{nuc}}+J(22)\right)\varphi_{2}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{1}=0.$
Eventually, we obtain two equations, where the kinetic part has the pattern $\frac{c_{i}^{2}}{2}\Delta-\varepsilon_{ii},$ that can be easily inverted. Indeed, we have
@@
(3.18) $\left(-\frac{c_{1}^{2}}{2}\Delta-\varepsilon_{11}\right)\delta\varphi_{1}+\left(1-\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}\right)\left(-\frac{c_{1}^{2}}{2}\Delta-\varepsilon_{11}\right)\varphi_{1}\\ +c_{1}^{2}\left(V_{\text{nuc}}+J(11)\right)\delta\varphi_{1}+2c_{1}^{2}J(1\delta 1)\varphi_{1}+c_{1}c_{2}J(\delta 12)\varphi_{2}+c_{1}c_{2}J(1\delta 2)\varphi_{2}+c_{1}c_{2}J(12)\delta\varphi_{2}-\varepsilon_{12}\delta\varphi_{2}\\ -\delta\varepsilon_{12}\varphi_{2}-\varepsilon_{12}\varphi_{2}-\left(\varepsilon_{11}\frac{\delta\theta\sin 2\theta}{c_{1}^{2}}+\delta\varepsilon_{11}\right)\varphi_{1}\\ +\left(c_{1}^{2}-\delta\theta\sin 2\theta\right)\left(V_{\text{nuc}}+J(11)\right)\varphi_{1}+(c_{1}c_{2}+\delta\theta\cos 2\theta)J(12)\varphi_{2}=0$
and